OpenStudy (anonymous):

How does the integral of (du/(u^2+4)) become 1/2 arctan (x+1/2) + C?

OpenStudy (anonymous):

Using Trig substituion

OpenStudy (anonymous):

u=tan(x) du=sec(x)^2 du

OpenStudy (anonymous):

How did u become over two, what happened to the four in the equation?

OpenStudy (anonymous):

substituting u=2tan(x) du=2sec(x)^2dx so u^2+4 is 4tan(x)^2+4 4(tan(x)^2 +1) \[1/4*2\int\limits \sec(x)^2/\sec(x)^2 d(x)\]